# Exact Enforcement of Temporal Continuity in Sequential Physics-Informed Neural Networks

### 15 Feb 2024 v1

@@ Abstract
The use of deep learning methods in scientific computing represents a potential paradigm shift in engineering problem solving. One of the most prominent developments is Physics-Informed Neural Networks (PINNs), in which neural networks are trained to satisfy partial differential equations (PDEs) and/or observed data. While this method shows promise, the standard version has been shown to struggle in accurately predicting the dynamic behavior of time-dependent problems. To address this challenge, methods have been proposed that decompose the time domain into multiple segments, employing a distinct neural network in each segment and directly incorporating continuity between them in the loss function of the minimization problem. In this work we introduce a method to exactly enforce continuity between successive time segments via a solution ansatz. This hard constrained sequential PINN (HCS-PINN) method is simple to implement and eliminates the need for any loss terms associated with temporal continuity. The method is tested for a number of benchmark problems involving both linear and non-linear PDEs. Examples include various first order time dependent problems in which traditional PINNs struggle, namely advection, Allen-Cahn, and Korteweg-de Vries equations. Furthermore, second and third order time-dependent problems are demonstrated via wave and Jerky dynamics examples, respectively. Notably, the Jerky dynamics problem is chaotic, making the problem especially sensitive to temporal accuracy. The numerical experiments conducted with the proposed method demonstrated superior convergence and accuracy over both traditional PINNs and the soft-constrained counterparts.
@@
Deepe learning, Physics-informed neural networks (PINNs), Temporal causality, Advection equation, Wave equation, Allen-Cahn equation, Jerk dynamics
journal: arXiv
@@ 1 Introduction
Physics-Informed Neural Networks, introduced by Raissi et al. raissi , encompass a general framework that can incorporate physical constraints via differential equations. In general, it involves using a deep neural network as a trial solution to a PDE-constrained optimization problem whose loss function includes residuals from governing equations, boundary and initial conditions, as well as observed data, if present. It is meshless, and the recent progress in deep learning tools has made its implementation through automatic differentiation a straightforward exercise. Indeed, it has been used in modeling a wide range of problems cai2021physics ; mahmoudabadbozchelou2022nn ; eivazi2022physics ; cai2021heat ; ghaffari2023deep ; niaki2021physics . Additionally, it shows promise for inverse problems as it can readily accommodate parameters and data into its structure tanyu2023deep ; yang2021b ; lu2021physics ; jagtap2022deep ; chen2020physics .
However, it has not seen widespread adoption in the computational science community. For one, it generally is not as accurate as traditional numerical methods such as the finite element and finite volume methods reddy2022finite . Additionally, the traditional PINN method is not robust as it often fails to converge due to the complexity of the loss landscape krishnapriyan2021characterizing . Moreover, time-dependent problems can suffer from a causality violation when the temporal variable is treated the same way as spatial variables wang2022respecting . Therefore, various improvements must be made before it can be an acceptable alternative for traditional numerical methods cuomo2022scientific . One common approach to simplify the process is to directly enforce initial and/or boundary conditions via specific solution ansatz or modified neural network architecture. This leaves only the PDE residuals (and possibly observable data) as terms in the loss function. For time-dependent problems, an additional approach is to divide the temporal domain into segments and use different neural networks for each time segment. This takes advantage of the inherent unidirectional nature of time and helps avoid spurious solutions by maintaining causality.
In this work, we propose a novel method for exactly imposing continuity between successive time segments, leading to hard constrained sequential PINNs (HCS-PINNs). We demonstrate its effectiveness for a set of carefully chosen time-dependent problems with increasing difficulty levels. The problem set encompasses both linear and non-linear equations involving first, second and third order derivatives in time. As demonstrated in the results, the HCS-PINN method can accurately predict the solution of time-dependent problems for which traditional PINNs struggle.
@@ 2 Background
where $\mathcal{F}$ is an operator which may consist of time and spatial derivatives. This is accompanied by initial and boundary conditions, respectively:
@@
$u(x,t)\approx f(x,t,\theta)=\arg\min_{\theta}\mathcal{L}({\theta}),$ (4)
@@ 3 Hard Constrained Sequential PINNs (HCS-PINNs)
Where $g(x,t)$ satisfies $\mathcal{I}[g(x,0)]=0$. However, this is a recursive form, where all previous neural networks must be evaluated. In order to avoid a recursive definition for the trial solution, we propose the following:
@@
Here $h_{N}$ and $h_{N+1}$ are auxiliary functions that obey certain conditions depending on the continuity of time derivatives. In order to break recursion and maintain $C^{0}$ continuity, we define $\tau=\frac{t-t_{N}}{t_{N+1}-t_{N}}$ and require the following:
@@
$C^{m}$ continuity can be obtained by additionally enforcing that the derivatives up to order $m$ vanish at $\tau=0$ and $\tau=1$. The auxiliary functions can be chosen by including polynomials from order 0 up to order $12m+$ 1 while enforcing these boundary conditions. Generally speaking, the continuity requirement between time segments is one less than the highest order time derivative in the problem, and choosing a value greater than this will unnecessarily over-constrain the problem. In this work, we will consider first, second, and third order problems with respect to time, which therefore leads to the choices for $h(\tau)$ listed in Table 1. Note that these functions satisfy a partition of unity, and can be interpreted as interpolation functions.
@@
The initial conditions for $M^{\text{th}}$-order problems (requiring $C^{M-1}$ continuity) are satisfied by taking the first $M$ terms of the following series:
@@ 4 Results
A framework was developed to test the accuracy and robustness of physics informed neural network algorithms for dynamic problems. The Python based Jax library jax2018github was used to develop the code, which employs feed forward neural networks. For all problems, boundary conditions are enforced directly in the solution ansatz or in the neural network architecture (see A). Therefore, the loss terms for HCS-PINNs solely consist of PDE residuals, whereas for SCS-PINNs, they include both initial and temporal continuity conditions. A dense distribution of random collocation points is used to sample each time segment, with mini-batching employed to speed up the training time, except for the Jerk equation where a full batch gradient descent is used. The optimization procedure starts with a fixed number of Adam kingma2014adam iterations. Then the L-BFGS liu1989limited iterations are performed until a convergence criterion is met or until it reaches a maximum number of iterations. Due to the stochastic nature of the mini-batching procedure, a simple five term moving average of the change in loss function from iteration to iteration is used as the convergence criterion. Additionally, a different (larger) mini-batch is used to evaluate the loss criterion to prevent false convergence. The activation function is the hyperbolic tangent ($tanh$), with all parameters initialized using the Glorot scheme glorot2010understanding . An empirical linear loss weighting scheme is used to further enforce temporal causality (see B). For any problem without an analytical solution, reference solutions were generated with the Chebfun package driscoll2014chebfun in Matlab. In each scenario, we analyze the relative $L^{2}$ error and training duration between HCS-PINN and SCS-PINN. The simulations were conducted using an Apple M1 Max processor, featuring a clock speed of 3.2 GHz.
We conduct numerical experiments by solving five dynamic problems with increasing level of complexity. For each problem, three scenarios are presented in which the temporal decomposition varies by the number of time steps ($nt$). The first two problems, the advection equation and wave equation, are linear and exhibit hyperbolic behavior. Despite their linear nature, the presence of a high characteristic velocity poses challenges for predicting solution over extended time intervals. The subsequent two problems, namely Allen-Cahn and Korteweg-de Vries (KdV) equations, are non-linear and commonly used in the PINN literature to benchmark algorithms. Lastly, we address a Jerk dynamics equation, known for its chaotic behavior.
@@ 4.1 Advection Equation
As the advection equation is first order in time, $C^{0}$ continuity needs to be enforced between adjacent time segments. When the fluid velocity is relatively large, vanilla PINNs struggle to capture the correct solution. In particular, PINN solutions often falsely converge to the zero solution, as seen in Figures 3 and 4. For SCS-PINN, taking 10 time segments is sufficient to avoid this for $c=50$, but not $c=100$. However, HCS-PINN captures the correct solution for all fluid speeds when taking 10 time steps, and captures up to $c=50$ when taking only 4.
Various other efforts have been made to mitigate this issue, including aforementioned temporal techniques as well as methods that encode characteristics into the architecture braga2022characteristics . Wang et al. wang2023expert suggested a range of techniques for obtaining optimal accuracy which include temporal and spatial Fourier feature embedding and grad norm weighting, among others. Indeed a relative $L^{2}$ error as low as O($10^{-4}$) was achieved for high transport velocity ($c=80$). In addition to the above techniques, they used a deep neural network (4 layers, 256 neurons in each layer, batch size of 8192) with a modified multi-layer perceptron (MLP). Here, we demonstrate that even with a relatively smaller network and batch size using a standard MLP, a relative $L^{2}$ error of O($10^{-3}$) can be reached for high advection velocity cases (i.e. $c=100$) with HCS-PINN.
@@ 4.2 Wave Equation
The second order derivative in time requires that we enforce $C^{1}$ continuity at time sub-domain interfaces. In addition to the temporal interpolation functions, the neural network is multiplied by the term $x(\pi-x)$ to strongly enforce zero Dirichlet conditions. For low wave speeds, such as $c=1$ or less, it is relatively easy to get accurate solutions. However, as the wave speed increases, this becomes increasingly difficult with vanilla PINNs. For example, when $c=10$, neither the HCS-PINN or SCS-PINN converges when only a single time segment is used. For the cases of 4 and 10 time windows, HCS-PINN gives less error than the corresponding SCS-PINN. Additionally, the 10 time step case actually takes less time than the 4 time step case, as it continually reaches the tolerance criterion earlier. These results are tabulated in Table 5.
@@ 4.3 Allen-Cahn Equation
$u_{t}-\lambda_{1}u_{xx}+\lambda_{2}u^{3}-\lambda_{2}u=0,$ (26)
@@
where $\lambda_{1}=0.0001$ and $\lambda_{2}=5$. The Allen-Cahn equation requires $C^{0}$ continuity between sequential neural networks. The interplay between the diffusion term and non-linear source terms makes it a particularly challenging problem, posing difficulties even for traditional numerical methods. The stiffness of the equation is dictated by the relative magnitude of the diffusion and reaction coefficients, i.e. $\lambda_{1}$ and $\lambda_{2}$. A number of methods, including Adaptive Time Sampling and bc-PINNs, were developed with this specific problem in mind.
@@ 4.4 Korteweg–de Vries (KdV) Equation
The solution is well captured by the SCS-PINN and especially the HCS-PINN within the time range [0,1], as shown in Figures 9 and 10. Again, the discontinuities at the time window interfaces in the error contour for SCS-PINN makes apparent the soft nature of continuity enforcement.
@@ 4.5 Chaotic Dynamics
Figure 11: Comparison with HCS-PINN and SCS-PINN solutions of Jerk dynamics equation with chebfun solution driscoll2014chebfun .
@@ 5 Summary
PINNs often struggle for time dependent problems due to issues with temporal causality. This has previously been addressed by performing a time-domain decomposition, employing a distinct neural network for each time segment, and ensuring temporal continuity through terms incorporated into the loss function. In this work we have introduced the HCS-PINN method, which eliminates the need to include these terms by strictly enforcing temporal continuity between successive neural networks.
Both methods were demonstrated on first, second, and third order problems in time. The advection and wave equations are linear, hyperbolic problems for which PINNs often erroneously converge to the zero solution if the specified characteristic speeds are large. HCS-PINN has shown to be a superior approach in avoiding this spurious solution. Allen-Cahn and Korteweg-de Vries are both nonlinear problems which are widely used as benchmarks in the literature. In all cases, the hard constrained PINNs outperformed the soft constrained counterparts in terms of accuracy. Finally, a jerky dynamics problem was presented that is chaotic in nature, making it very sensitive to initial conditions. HCS-PINN was able to obtain the correct solution with only 4 time steps while SCS-PINN was not, and for 10 time steps its accuracy was better by an order of magnitude.
@@ Appendix B Empirical Causal Weighting
With the exception of the Jerk dynamics problem, $t_{\text{max}}$ is taken to be the maximum time in the problem, not in the time step. For all problems, $C_{T}=10$ was used as causal weighting parameter.

### 07 Mar 2024 v2

@@ Abstract
The use of deep learning methods in scientific computing represents a potential paradigm shift in engineering problem solving. One of the most prominent developments is Physics-Informed Neural Networks (PINNs), in which neural networks are trained to satisfy partial differential equations (PDEs). While this method shows promise, the standard version has been shown to struggle in accurately predicting the dynamic behavior of time-dependent problems. To address this challenge, methods have been proposed that decompose the time domain into multiple segments, employing a distinct neural network in each segment and directly incorporating continuity between them in the loss function of the minimization problem. In this work we introduce a method to exactly enforce continuity between successive time segments via a solution ansatz. This hard constrained sequential PINN (HCS-PINN) method is simple to implement and eliminates the need for any loss terms associated with temporal continuity. The method is tested for a number of benchmark problems involving both linear and non-linear PDEs. Examples include various first order time dependent problems in which traditional PINNs struggle, namely advection, Allen-Cahn, and Korteweg-de Vries equations. Furthermore, second and third order time-dependent problems are demonstrated via wave and Jerky dynamics examples, respectively. Notably, the Jerky dynamics problem is chaotic, making the problem especially sensitive to temporal accuracy. The numerical experiments conducted with the proposed method demonstrated superior convergence and accuracy over both traditional PINNs and the soft-constrained counterparts.
@@
Deep learning, Physics-informed neural networks (PINNs), Causality, Temporal continuity, Chaotic equations
journal: Journal of Computational Physics
@@ 1 Introduction
Physics-Informed Neural Networks, introduced by Raissi et al. raissi , encompass a general framework that can incorporate physical constraints via differential equations. In general, it involves using a deep neural network as a trial solution to a PDE-constrained optimization problem whose loss function includes residuals from governing equations, boundary and initial conditions, as well as observed data, if present. It is meshless, and the recent progress in deep learning tools has made its implementation through automatic differentiation a straightforward exercise. Indeed, it has been used in modeling a variety of problems, ranging from fluid mechanics cai2021physics ; mahmoudabadbozchelou2022nn ; eivazi2022physics ; biswas2023three , heat transfer cai2021heat , solid mechanics ghaffari2023deep to subsurface systems sarma2023variational and chemical processes niaki2021physics . Additionally, it shows promise for inverse problems as it can readily accommodate parameters and data into its structure tanyu2023deep ; yang2021b ; lu2021physics ; jagtap2022deep ; chen2020physics ; serebrennikova2024physics .
However, it has not seen widespread adoption in the computational science community. For one, it generally is not as accurate as traditional numerical methods such as the finite element and finite volume methods reddy2022finite . Additionally, the standard PINN method is not robust as it often fails to converge due to the complexity of the loss landscape krishnapriyan2021characterizing . Moreover, time-dependent problems can suffer from a causality violation when the temporal variable is treated the same way as spatial variables wang2022respecting . Therefore, various improvements must be made before it can be an acceptable alternative for traditional numerical methods cuomo2022scientific . One common approach to simplify the process is to directly enforce initial and/or boundary conditions via specific solution ansatz or modified neural network architecture. This leaves only the PDE residuals (and possibly observable data) as terms in the loss function. For time-dependent problems, an additional approach is to divide the temporal domain into segments and use different neural networks for each one. This takes advantage of the inherent unidirectional nature of time and helps avoid spurious solutions by maintaining causality.
In this work, we propose a novel method for exactly imposing continuity between successive time segments, leading to hard constrained sequential PINNs (HCS-PINNs). We demonstrate its effectiveness for a set of carefully chosen time-dependent problems with increasing difficulty. The problem set encompasses both linear and non-linear equations involving first, second and third order derivatives in time. As demonstrated in the results, the HCS-PINN method can accurately predict the solution of time-dependent problems for which traditional PINNs struggle.
@@ 2 Background
where $\mathcal{F}$ is an operator which may consist of time and spatial derivatives. This is accompanied by
@@
$u(x,t)\approx f(x,t,\theta);\quad\theta=\arg\min_{\theta^{*}}\mathcal{L}(\theta^{*}),$ (4)
@@ 3 Hard Constrained Sequential PINNs (HCS-PINNs)
Where $g(x,t)$ satisfies $\mathcal{I}[g(x,0)]=0$. However, this is a recursive form, where neural network solutions for all previous time windows must be evaluated to compute the solution for the current time window. In order to avoid a recursive definition for the trial solution, we propose the following:
@@
Here, $h_{N}$ and $h_{N+1}$ are auxiliary functions chosen to obey certain conditions depending on the desired continuity of time derivatives. In order to break recursion and maintain $C^{0}$ continuity, we define $\tau=\frac{t-t_{N}}{t_{N+1}-t_{N}}$ and require the following:
@@
$C^{m}$ continuity can be obtained by additionally enforcing that the derivatives up to order $m$ vanish at $\tau=0$ and $\tau=1$. The auxiliary functions can be chosen by including polynomials from order 0 up to order $12m+$ 1 while enforcing these boundary conditions. Generally speaking, the continuity requirement between time segments is one less than the highest order time derivative in the problem, and choosing a value greater than this will unnecessarily over-constrain the problem. In this work, we will consider first, second, and third order problems with respect to time, which therefore leads to the choices for $h(\tau)$ listed in Table 1. Note that these functions satisfy a partition of unity, and can be interpreted as interpolation functions (see Figure 3).
Figure 3: Interpolation functions for $C^{0}$, $C^{1}$, and $C^{2}$ continuity.
@@
The initial conditions for $M^{\text{th}}$-order problems in time (requiring $C^{M-1}$ continuity) are satisfied by taking the first $M$ terms of the following series:
@@ 4 Results
A framework was developed to test the accuracy and robustness of physics informed neural network algorithms for dynamic problems. The Python based Jax library jax2018github was used to develop the code, which employs feed forward neural networks. For all problems, boundary conditions are enforced directly in the solution ansatz or encoded in the neural network architecture (see A). Therefore, the loss terms for HCS-PINNs solely consist of PDE residuals, whereas for SCS-PINNs, they include both initial and temporal continuity conditions. A dense distribution of random collocation points is used to sample each time segment, with mini-batching employed to speed up the training time, except for the Jerk equation where a full batch gradient descent is used. The optimization procedure starts with a fixed number of Adam kingma2014adam iterations. Then the L-BFGS liu1989limited iterations are performed until a convergence criterion is met or until it reaches a maximum number of iterations. Due to the stochastic nature of the mini-batching procedure, a simple five term moving average of the change in loss function from iteration to iteration is used as the convergence criterion. Additionally, a different (larger) mini-batch is used to evaluate the loss criterion to prevent false convergence. The activation function is the hyperbolic tangent ($tanh$), with all parameters initialized using the Glorot scheme glorot2010understanding . An empirical linear loss weighting scheme is used to further enforce temporal causality (see B). For any problem without an analytical solution, reference solutions were generated with the Chebfun package driscoll2014chebfun in Matlab. In each scenario, we analyze the relative $L^{2}$ error and training duration between HCS-PINN and SCS-PINN. The simulations were conducted using an Apple M1 Max processor, featuring a clock speed of 3.2 GHz and a memory of 32 GB.
We conduct numerical experiments by solving five dynamic problems with increasing level of complexity. For each problem, three scenarios are presented in which the temporal decomposition varies by the number of time steps ($nt$). The first two problems, the advection equation and wave equation, are linear and exhibit hyperbolic behavior. Despite their linear nature, the presence of a high characteristic velocity poses challenges for predicting solution over extended time intervals. The subsequent two problems, namely Allen-Cahn and Korteweg-de Vries (KdV) equations, are non-linear and commonly used in the PINN literature to benchmark algorithms. Lastly, we address a Jerky dynamics equation, known for its chaotic behavior.
@@ 4.1 Advection Equation
As the advection equation is first order in time, $C^{0}$ continuity needs to be enforced between adjacent time segments. In our implementation, the PDE has been normalized by the fluid velocity. We found that this normalization sufficiently enforced the initial condition for SCS-PINNs. Alternatively, the penalty parameter $\lambda_{I}$ could be increased to achieve the same effect.
When the fluid velocity is relatively large, vanilla PINNs struggle to capture the correct solution. In particular, PINN solutions often falsely converge to the zero solution, as seen in Figures 4 and 5. For SCS-PINN, taking 10 time segments is sufficient to avoid this for $c=50$, but not for $c=100$. However, HCS-PINN captures the correct solution for all fluid speeds when taking 10 time steps, and captures up to $c=50$ when taking only 4.
Various other efforts have been made to mitigate this issue, including aforementioned temporal techniques as well as methods that encode characteristics into the architecture braga2022characteristics . Wang et al. wang2023expert suggested a range of techniques for obtaining optimal accuracy which include temporal and spatial Fourier feature embedding and grad norm weighting, among others. Indeed a relative $L^{2}$ error as low as O($10^{-4}$) was achieved for high transport velocity ($c=80$). In addition to the above techniques, they used a deep neural network (4 layers, 256 neurons in each layer, batch size of 8192) with a modified multi-layer perceptron (MLP). Here, we demonstrate that even with a relatively smaller network and a batch size (4 layers, 32 neurons in each layer, batch size of 128) using a standard MLP, a relative $L^{2}$ error of O($10^{-3}$) can be reached for high advection velocity cases (i.e. $c=100$) with HCS-PINN.
@@ 4.2 Wave Equation
The second order derivative in time requires that we enforce $C^{1}$ continuity at time sub-domain interfaces. In addition to the temporal interpolation functions, the neural network is multiplied by the term $x(\pi-x)$ to strongly enforce zero Dirichlet boundary conditions. Similar to the advection equation, the PDE was scaled by the square of the wave speed in lieu of altering the weighting parameter $\lambda_{I}$ for SCS-PINNs. For low wave speeds, such as $c=1$ or less, it is relatively easy to get accurate solutions (see Figure 6). However, as the wave speed increases, this becomes increasingly difficult with vanilla PINNs. For example, when $c=10$, neither the HCS-PINN or SCS-PINN converges when only a single time segment is used, as shown in Figure 7. For the cases of 4 and 10 time windows, HCS-PINN gives less error than the corresponding SCS-PINN. Additionally, the 10 time step case actually takes less time than the 4 time step case, as it continually reaches the tolerance criterion earlier. These results are tabulated in Table 5.
@@ 4.3 Allen-Cahn Equation
$u_{t}-\lambda_{1}u_{xx}+\lambda_{2}u^{3}-\lambda_{2}u=0,\>x\in[-1,1],\>t\in[0,1]$ (26)
@@
where $\lambda_{1}=0.0001$ and $\lambda_{2}=5$. The Allen-Cahn equation requires $C^{0}$ continuity between sequential neural networks. The interplay between the diffusion term and non-linear source terms makes it a particularly challenging problem, posing difficulties even for traditional numerical methods. The stiffness of the equation is dictated by the relative magnitude of the diffusion and reaction coefficients, i.e. $\lambda_{1}$ and $\lambda_{2}$, respectively. A number of methods, including Adaptive Time Sampling and bc-PINNs, were developed with this specific problem in mind.
@@ 4.4 Korteweg–de Vries (KdV) Equation
The solution is well captured by the SCS-PINN and especially the HCS-PINN within the time range [0,1], as shown in Figures 10 and 11. Again, the discontinuities at the time window interfaces in the error contour for SCS-PINN makes apparent the soft nature of continuity enforcement.
@@ 4.5 Chaotic Dynamics
Figure 12: Comparison with HCS-PINN and SCS-PINN solutions of Jerky dynamics equation with chebfun solution driscoll2014chebfun .
@@ 5 Summary
PINNs often struggle for time-dependent problems due to issues with temporal causality. This has previously been addressed by performing a time-domain decomposition, employing a distinct neural network for each time segment, and ensuring temporal continuity through terms incorporated into the loss function. In this work we have introduced the HCS-PINN method, which eliminates the need to include these terms by strictly enforcing temporal continuity between successive neural networks.
Both methods were demonstrated on first, second, and third order problems in time. The advection and wave equations are linear, hyperbolic problems for which PINNs often erroneously converge to the zero solution if the specified characteristic speeds are large. HCS-PINN has shown to be a superior approach in avoiding this spurious solution. Allen-Cahn and Korteweg-de Vries are both nonlinear problems which are widely used as benchmarks in the literature. In all cases, the hard constrained PINNs outperformed the soft constrained counterparts in terms of accuracy. Finally, a jerky dynamics problem was presented that is chaotic in nature, making it very sensitive to initial conditions. HCS-PINN was able to obtain the correct solution with only 5 time steps while SCS-PINN was not, and for 10 time steps its accuracy was better by an order of magnitude.
@@ Appendix B Empirical Causal Weighting
With the exception of the Jerky dynamics problem, $t_{\text{max}}$ is taken to be the maximum time in the problem, not in the time step. For all problems, $C_{T}=10$ was used as causal weighting parameter.
@@ Appendix C Hyper-Parameters
A list of hyper-parameters used in various problems is found in Table 9.